Hello! I've tried to approach this from a substiution standpoint but I'm at a loss :
Find the number of solutions for the system
(the following equations are all together in one bracket)
{f + g + h = 4
f = 5 - h - g
g + h = 7 - f
If you could help me on how I should approach this problem I would be very grateful!! Thanks!!!!!

Answer by venugopalramana(3286)

Hello! I've tried to approach this from a substiution standpoint but I'm at a loss :
Find the number of solutions for the system
(the following equations are all together in one bracket)
{f + g + h = 4.........................................I
f = 5 - h - g.....OR.....F+G+H=5.....................II
g + h = 7 - f.....OR.....F+G+H=7
If you could help me on how I should approach this problem I would be very grateful!! Thanks!!!!!
THESE 3 ARE IDENTICAL ON LHS BUT HAVE DIFFERENT VALUES ON RHS.HENCE THEY ARTE INCONSISTENT AND NO SOLUTION EXISTS WHICH CAN SATISFY ALL 3 EQNS.